# Quantum Features of the Thermal Two-Qubit Quantum Rabi Model in Ultra- and Deep-Strong Regimes

### v1

@@ Abstract
Quantum correlations and nonclassical states are indispensable resources for advancing quantum technologies, and their resilience at finite temperatures is crucial for practical experimental implementations. The two-qubit quantum Rabi model (2QQRM), a natural extension of the quantum Rabi model, describes two qubits coupled to a single bosonic mode and has been extensively studied in cavity quantum electrodynamics, superconducting circuits, and quantum information science. In this work, we investigate the persistence of quantum correlations and nonclassical states in the 2QQRM at thermal equilibrium, focusing on the ultrastrong and deep strong coupling regimes. Through a systematic analysis of quantumness quantifiers, we demonstrate the emergence of long-lived quantum correlations, even in the presence of thermal noise. Notably, we uncover striking phenomena arising from the interplay between detuning and deep strong-coupling: in the high-frequency limit, where the qubit energy exceeds the cavity-mode energy, quantum criticality emerges, leading to a high degree of photon squeezing. In contrast, the opposite regime is characterized by robust qubit-qubit quantum correlations. Importantly, we show that both dispersive regimes exhibit quantum features that are remarkably robust to parameter fluctuations, making them advantageous for maintaining quantum coherence. These results highlight the exceptional resilience of quantum resources in the 2QQRM and provide valuable insights for developing quantum technologies operating under realistic, finite-temperature conditions.
@@ I Introduction
The quantum Rabi model (QRM), describing the interaction between a two-level atom (qubit) and a single-mode cavity field, serves as the fundamental framework for studying light-matter interactions [42, 43, 44, 45]. This model has been widely studied across various fields, such as quantum optics [46], entanglement generation [47], quantum thermodynamics [48, 49, 50, 51, 52], and the study of quantum phase transitions (QPTs) [53, 54, 55, 56, 57, 58]. Modifications to the QRM, such as the anisotropic quantum Rabi model (AQRM) [59, 60] and the Jaynes-Cummings model (JCM) [61], provide even richer dynamics by incorporating different coupling terms and interactions. Both the AQRM and the JCM are known to exhibit first-order quantum phase transitions, which influence quantum thermalization and the persistence of quantum correlations at thermal equilibrium [62, 63, 64, 65].
@@ II The Model
To conclude our review, we present in Fig. 1 the energy spectrum (with blue and red indicating different parity states) and steady-state population $P_{0}$ across different temperatures for degenerate qubits in various regimes. Figs. 1(b) and (c) show the resonant interaction regime ($\omega=\Delta$), where the energy levels exhibit significant anharmonicity, transitioning to a degenerate harmonic spectrum in the deep strong coupling limit. Figs. 1(d) and (e) illustrate the high-frequency qubit regime ($\Delta\gg\omega,g$), revealing distinct behaviors before and after the critical coupling $g_{c}=\frac{\sqrt{\omega\Delta}}{2}$. Figs. 1(f) and (g) show the adiabatic quantum oscillator regime ($\omega\gg\Delta,g$), where the system evolves adiabatically and the qubits track the oscillator’s evolution in a nearly flat potential.
Lastly, numerical computations throughout this work were carried out using QuTiP [101], with a variable Fock basis employed to achieve convergence.
@@ III.2 Quadrature Squeezing
It is worth noting that previous studies, such as Refs. [107, 108], have explored the squeezing of output quadratures in the USC and DSC regimes using dressed light-matter jump operators for a vacuum input field. Interestingly, these studies found that for any open system in its ground state, output squeezing cannot be produced, even if the ground state itself is squeezed. Since most of the squeezing in the two-qubit Rabi model is present in the ground state, this generalized measure leads to a zero degree of squeezing in the thermal state when the light-matter interaction is active. For this reason, we focus on evaluating the degree of squeezing after turning off the light-matter interaction. This process effectively transforms virtual photons into real, measurable photons, revealing a clearer signature of squeezing in the system. Thus, after switching off the interaction, we capture the squeezing behavior by computing $\zeta^{2}$, highlighting the nonclassical features of the state.
@@ IV Results and discussion
Consistent with the observations in Fig. 5, the quantum properties of the qubits remain nearly constant at low temperatures ($\textbf{T}\lesssim 0.1$), regardless of the chosen parameters. However, at higher temperatures, the behavior of different quantities diverges. Specifically, concurrence (Fig. 9(a)) and quantum discord (Fig. 9(c)) decay to zero, following a trend similar to the squeezing parameter and second-order correlation function. While negativity (Fig. 9(b)) and local quantum uncertainty (LQU, Fig. 9(d)) also decrease significantly, they remain measurable for certain parameter sets, even at $\textbf{T}=1$. In contrast, mutual information (Fig. 9(f)) generally vanishes rapidly with increasing temperature across most parameter configurations, with the exception of the resonant case where $g=\omega=1$, where it survives up to $\textbf{T}=5$. Notably, at lower temperatures, some parameter sets (e.g., $g=\omega=1$) yield nonzero mutual information while simultaneously exhibiting zero quantum discord, indicating that the correlations are predominantly classical in this regime. Conversely, parameter sets that produce both significant mutual information and quantum discord (e.g., $g=6$ and $\omega=60$) confirm the persistence of quantum correlations within the system. Lastly, coherence (Fig. 9(e)) initially increases with temperature before vanishing at higher T, which is in agreement with recent studies [87]. This transient enhancement reveals a counterintuitive effect where thermal excitations temporarily amplify coherence, only for it to eventually degrade as the temperature continues to rise, thereby diminishing the system’s quantum properties.
@@ V CONCLUDING REMARKS
The 2QQRM retains key nonclassical features of the QRM – particularly photon antibunching, though now peaking at half the characteristic coupling strength – while exhibiting several novel quantum behaviors. Although squeezing has been observed in the QRM in the high-frequency regime away from criticality at zero temperature [38] or under resonant conditions in thermal equilibrium [65], we demonstrate that the 2QQRM displays markedly different behavior: strong squeezing emerges near criticality but becomes completely suppressed in the superradiant phase. Most importantly, we uncover three distinctive phenomena in the 2QQRM that represent either qualitative departures from or entirely new features compared to the single-qubit case: (i) diverging thermalization timescales; (ii) persistent entanglement between qubits (a feature fundamentally impossible in the single-qubit QRM); and (iii) robust quantum coherence with finite local quantum uncertainty – none of which have been systematically investigated for the QRM under thermal equilibrium conditions to date.
@@
Our theoretical proposal to achieve nonclassical equilibrium states is both fundamentally significant and experimentally feasible, as recent advances in strong coupling and low-temperature techniques demonstrate. The USC regime, characterized by $g/\omega>0.1$, has been experimentally realized for more than a decade [28] and continues to be realized in several systems [30]. The DSC regime, with $g/\omega\approx 1.34$ and a temperature of approximately 45 mK [31], is now within reach, with thermal entanglement observed in the quantum Rabi model. Deep-strong coupling considering the QRM is also observed with a single-trapped ion and, more recently, with trapped atoms [95], remarkably achieving couplings of $\sim 6.5$ greater than the field frequency [123]. Additionally, novel fluxonium-based superconducting devices are being developed to detect virtual photons in the USC regime [124]. Hybrid semiconductor-superconductor platforms further enhance flexibility in the implementation of the QRM [125], underscoring the exciting potential to realize nonclassical equilibrium states within these coupling regimes. Finally, USC is also reached using cold atoms [126] and metamaterials [127].

### v2

@@ Abstract
Quantum correlations and non-classical states are indispensable resources for advancing quantum technologies, and their resilience at finite temperatures is crucial for practical experimental implementations. The two-qubit quantum Rabi model (2QQRM), a natural extension of the quantum Rabi model, describes two qubits coupled to a single bosonic mode and has been extensively studied in cavity quantum electrodynamics, superconducting circuits, and quantum information science. In this work, we investigate the persistence of quantum correlations and non-classical states in the 2QQRM at thermal equilibrium, focusing on the ultrastrong and deep strong coupling regimes. Through a systematic analysis of quantumness quantifiers, we demonstrate the emergence of long-lived quantum correlations, even in the presence of thermal noise. Notably, we uncover striking phenomena arising from the interplay between detuning and deep strong-coupling: in the high-frequency limit, where the qubit energy exceeds the cavity-mode energy, quantum criticality emerges, leading to a high degree of photon squeezing. In contrast, the opposite regime is characterized by robust qubit-qubit quantum correlations. Importantly, we show that both dispersive regimes exhibit quantum features that are remarkably robust to parameter fluctuations, making them advantageous for maintaining quantum coherence. These results highlight the exceptional resilience of quantum resources in the 2QQRM and provide valuable insights for developing quantum technologies operating under realistic, finite-temperature conditions.
@@ I Introduction
The quantum Rabi model (QRM), describing the interaction between a two-level atom (qubit) and a single-mode cavity field, serves as the fundamental framework for studying light-matter interactions [42, 43, 44, 45]. This model has been widely studied across various fields, such as quantum optics [46], entanglement generation [47], quantum thermodynamics [48, 49, 50, 51, 52], and the study of quantum phase transitions (QPTs) [53, 54, 55, 56, 57, 58, 59]. Modifications to the QRM, such as the anisotropic quantum Rabi model (AQRM) [60, 61] and the Jaynes-Cummings model (JCM) [62], provide even richer dynamics by incorporating different coupling terms and interactions. Both the AQRM and the JCM are known to exhibit first-order quantum phase transitions, which influence quantum thermalization and the persistence of quantum correlations at thermal equilibrium [63, 64, 65, 66].
@@ II The Model
Figure 2: Regimes of the two-qubit quantum Rabi model. Key features: (1) RWA and non-RWA dispersive limits show fundamentally different entanglement structures, (2) deep strong coupling reveals displaced Fock product states, (3) the adiabatic regime maintains qubit correlations despite a harmonic spectrum, and (4) the high-frequency limit exhibits squeezing and superradiant-like behavior.
@@
Summarizing the theoretical results for limit cases aforementioned, we elaborate the diagram presented in Fig. 2, where we show in a concise way the differences and specifics of each parameter regime. In addition, numerical computations throughout this work were carried out using QuTiP [102]. In all simulations, the full Hamiltonian of Eq. (1) was considered. To ensure accurate results without expending unnecessary computational labor, we use a variable Fock basis, checking its precision in each case by computing the bosonic commutator of the Fock annihilation and creation operators, which presents an error smaller than $10^{-4}$. Besides, by considering larger bases during preliminary tests, we verify the convergence of the results for all the quantities studied throughout this work.
To conclude, we present in Fig. 1 the energy spectrum (with blue and red indicating different parity states) and steady-state population $P_{0}$ across different temperatures for degenerate qubits in various regimes. Figs. 1(b) and (c) show the resonant interaction regime ($\omega=\Delta$), where the energy levels exhibit significant anharmonicity, transitioning to a degenerate harmonic spectrum in the deep strong coupling limit. Figs. 1(d) and (e) illustrate the high-frequency qubit regime ($\Delta\gg\omega,g$), revealing distinct behaviors before and after the critical coupling $g_{c}=\frac{\sqrt{\omega\Delta}}{2}$. Figs. 1(f) and (g) show the adiabatic quantum oscillator regime ($\omega\gg\Delta,g$), where the system evolves adiabatically and the qubits track the oscillator’s evolution in a nearly flat potential.
@@ III.2 Quadrature Squeezing
It is worth noting that previous studies, such as Refs. [108, 109], have explored the squeezing of output quadratures in the USC and DSC regimes using dressed light-matter jump operators for a vacuum input field. Interestingly, these studies found that for any open system in its ground state, output squeezing cannot be produced, even if the ground state itself is squeezed. Since most of the squeezing in the two-qubit Rabi model is present in the ground state, this generalized measure leads to a zero degree of squeezing in the thermal state when the light-matter interaction is active. For this reason, we focus on evaluating the degree of squeezing after turning off the light-matter interaction. This process effectively transforms virtual photons into real, measurable photons, revealing a clearer signature of squeezing in the system. Thus, after switching off the interaction, we capture the squeezing behavior by computing $\zeta^{2}$, highlighting the nonclassical features of the state. To effectively convert virtual excitations into real photons, the light-matter coupling must be switched non-adiabatically, on timescales shorter than the inverse of the system’s characteristic frequencies [107]. Currently, switching times of a few nanoseconds – corresponding to approximately 10101010100100100100 cycles of GHz frequencies – are achievable [110]. While finite switching durations and decoherence could, in principle, introduce imperfections, the available experimental techniques are adequate for observing the predicted photonic effects.
@@ IV Results and discussion
Consistent with the observations in Fig. 6, the quantum properties of the qubits remain nearly constant at low temperatures ($\textbf{T}\lesssim 0.1$), regardless of the chosen parameters. However, at higher temperatures, the behavior of different quantities diverges. Specifically, concurrence (Fig. 10(a)) and quantum discord (Fig. 10(c)) decay to zero, following a trend similar to the squeezing parameter and second-order correlation function. While negativity (Fig. 10(b)) and local quantum uncertainty (LQU, Fig. 10(d)) also decrease significantly, they remain measurable for certain parameter sets, even at $\textbf{T}=1$. In contrast, mutual information (Fig. 10(f)) generally vanishes rapidly with increasing temperature across most parameter configurations, with the exception of the resonant case where $g=\omega=1$, where it survives up to $\textbf{T}=5$. Notably, at lower temperatures, some parameter sets (e.g., $g=\omega=1$) yield nonzero mutual information while simultaneously exhibiting zero quantum discord, indicating that the correlations are predominantly classical in this regime. Conversely, parameter sets that produce both significant mutual information and quantum discord (e.g., $g=6$ and $\omega=60$) confirm the persistence of quantum correlations within the system.
Lastly, coherence (Fig.10(e)) displays remarkable resilience to thermal effects. Interestingly, it initially increases with temperature before gradually diminishing at higher values of T. This counterintuitive behavior aligns qualitatively with recent findings[88], although their analysis considered thermal equilibrium only for the field mode, not the qubits. The observed enhancement reveals a regime where moderate thermal excitations can amplify coherence—a trend similarly seen in mutual information (MI) and local quantum uncertainty (LQU). While conventional nonclassicality measures such as entanglement and squeezing typically decay rapidly and exhibit strong quantum features only at zero temperature, these correlation-based quantifiers persist strikingly well. Beyond the optimal temperature point, coherence, MI, and LQU degrade much more slowly than other quantum resources across a wide range of parameters. This distinctive ”thermal persistence” effect could offer significant advantages for applications demanding robustness against thermal noise.
Table 1: Comparative analysis of nonclassicality metrics between the two-qubit quantum Rabi model (2QQRM) and the standard QRM.
Metric New Insights from the 2QQRM
$G^{(2)}(0)$ Reveals stronger antibunching at lower coupling strengths compared to QRM. The most nonclassical photon statistics emerge at approximately half the coupling strength of the QRM. At higher coupling, the statistics approach those of coherent states. Detuning serves as a powerful control parameter – its role was underexplored in prior QRM studies.
Squeezing Observed beyond the resonant case, especially near the critical point in the USC regime. Unlike earlier QRM analyses limited to resonance, the 2QQRM unveils previously unreported squeezing behavior in off-resonant and critical regimes.
Negativity Pronounced near resonance in the USC regime, but decreases when moving into the dispersive or DSC regime. Reflects strong but shifting qubit-field entanglement as coupling increases. The behavior is qualitatively similar to QRM.
Concurrence Qubit-qubit entanglement is absent in the single-qubit QRM but emerges naturally in 2QQRM due to indirect interaction via the shared field. Especially prominent in the DSC regime under the high-frequency qubit condition and robust at moderate temperatures.
Quantum Discord Tracks with concurrence, showing robust nonclassical correlations across broad parameter ranges. Useful when entanglement vanishes but quantum correlations persist.
Local Quantum Uncertainty (LQU) Exhibits high values near resonance and in the USC regime, with detuning amplifying its magnitude in the transition to DSC. While prior QRM studies focused on nonequilibrium or closed scenarios, this work explores its thermal equilibrium behavior.
Quantum Coherence Coherence unexpectedly increases at low temperatures, contrasting with typical decoherence behavior in QRM. Highlights the interplay between thermal effects and effective qubit-qubit interactions in the 2QQRM.
Table 2: Experimental platforms in the USC/DSC regime.
Platform (Ref.) $\omega/2\pi$ $g/\omega$ $T$ $k_{B}T/\hbar\omega$ Switching
Supercond. [125] 4.5 GHz 1.34 10 mK 0.046 1–100 ns
Semiconductor [30] 363 THz 143 10 mK $9.1\!\times\!10^{-7}$ 1–100 ps
Trapped ions [126] 1 MHz 650 1 mK 3.3 1–10 $\mu$s
Ultracold atoms [127] 100 Hz 100 1 $\mu$K 3.3 1–100 $\mu$s
Cavity QED [128] 5 GHz 0.5 15 mK 0.066 0.1–10 ns
@@ V CONCLUDING REMARKS
The 2QQRM retains key nonclassical features of the QRM – particularly photon antibunching, though now peaking at half the characteristic coupling strength – while exhibiting several novel quantum behaviors. Although squeezing has been observed in the QRM in the high-frequency regime away from criticality at zero temperature [38] or under resonant conditions in thermal equilibrium [66], we demonstrate that the 2QQRM displays markedly different behavior: strong squeezing emerges near criticality but becomes completely suppressed in the superradiant phase. Most importantly, we uncover three distinctive phenomena in the 2QQRM that represent either qualitative departures from or entirely new features compared to the single-qubit case: (i) diverging thermalization timescales; (ii) persistent entanglement between qubits (a feature fundamentally impossible in the single-qubit QRM); and (iii) robust quantum coherence with finite local quantum uncertainty – none of which have been systematically investigated for the QRM under thermal equilibrium conditions to date. For clarity, we compile in Table 1 a comparative summary of the nonclassicality measures explored in this study. This table distills the main findings into an accessible format, highlighting both shared and distinctive features of the two models.
@@
Our theoretical proposal to achieve nonclassical equilibrium states is both fundamentally significant and experimentally feasible, as recent advances in strong coupling and low-temperature techniques demonstrate. The USC regime, characterized by $g/\omega>0.1$, has been experimentally realized for more than a decade [28] and continues to be realized in several systems [30]. The DSC regime, with $g/\omega\approx 1.34$ and a temperature of approximately 45 mK [31], is now within reach, with thermal entanglement observed in the quantum Rabi model. Crucially, these implementations also permit active control of the light-matter coupling: superconducting circuits achieve ns-scale switching via magnetic flux tuning, while semiconductor systems enable ultrafast ps-scale control through optical pulses.
Deep-strong coupling considering the QRM is also observed with a single-trapped ion and, more recently, with trapped atoms [96], remarkably achieving couplings of $\sim 6.5$ greater than the field frequency [126]. In these atomic systems, the coupling can be modulated on $\mu$s timescales through trap potential adjustments or laser control. Additionally, novel fluxonium-based superconducting devices are being developed to detect virtual photons in the USC regime [129], with the added capability of in situ coupling strength modulation at nanosecond speeds. Hybrid semiconductor-superconductor platforms further enhance flexibility in the implementation of the QRM [130], combining fast switching with robust coherence properties.
The switching capabilities across platforms create new opportunities for dynamical control of quantum effects: from ultrafast (ps-scale) modulation in semiconductor systems ideal for studying non-equilibrium phenomena, to the more gradual ($\mu$s-scale) control in atomic systems suitable for observing thermalization processes. This underscores the exciting potential to realize nonclassical equilibrium states within these coupling regimes. Finally, USC is also reached using cold atoms [127] and metamaterials [131], each offering distinct approaches to coupling control – from optical manipulation in atomic systems to microwave tuning in metamaterials.
To contextualize these theoretical results and demonstrate the feasibility of implementation, we provide in Table 2 an overview of key experimental platforms that have reached the ultrastrong and deep-strong coupling regimes, including their characteristic switching timescales. The table includes details on relevant parameters such as coupling ratios, operating temperatures, and quantum system types, thereby serving as a practical reference for identifying systems where the predicted nonclassical equilibrium states of the 2QQRM may be observed. This comparison highlights the growing experimental capacity to probe strong light-matter interactions under thermal equilibrium, bridging theoretical proposals with cutting-edge technologies.